# Harmony in Divergence: Towards Fast, Accurate, and Memory-efficient Zeroth-order LLM Fine-tuning

### v1

@@ Abstract
Large language models (LLMs) excel across various tasks, but standard first-order (FO) fine-tuning demands considerable memory, significantly limiting real-world deployment. Recently, zeroth-order (ZO) optimization stood out as a promising memory-efficient training paradigm, avoiding backward passes and relying solely on forward passes for gradient estimation, making it attractive for resource-constrained scenarios. However, ZO method lags far behind FO method in both convergence speed and accuracy. To bridge the gap, we introduce a novel layer-wise divergence analysis that uncovers the distinct update pattern of FO and ZO optimization. Aiming to resemble the learning capacity of FO method from the findings, we propose Divergence-driven Zeroth-Order (DiZO) optimization. DiZO conducts divergence-driven layer adaptation by incorporating projections to ZO updates, generating diverse-magnitude updates precisely scaled to layer-wise individual optimization needs. Our results demonstrate that DiZO significantly reduces the needed iterations for convergence without sacrificing throughput, cutting training GPU hours by up to 48% on various datasets. Moreover, DiZO consistently outperforms the representative ZO baselines in fine-tuning RoBERTa-large, OPT-series, and Llama-series on downstream tasks and, in some cases, even surpasses memory-intensive FO fine-tuning.
Machine Learning, ICML
@@ 1 Introduction
Fine-tuning pre-trained large language models (LLMs) with backpropagation demonstrates superior performance for many natural language processing tasks (Yang et al., 2019; Liu et al., 2019; Talmor et al., 2018; Chowdhery et al., 2023; Zheng et al., 2020). However, the extensive parameterization imposes a substantial memory burden, limiting their practicality for widespread downstream applications. In line with the neural scaling laws (Hoffmann et al., 2022; Kaplan et al., 2020), next-generation LLMs continue to increase in parameter count. Specifically, model sizes are expanding at a rate of 410× every two years, dramatically outpacing the scaling of DRAM bandwidth (1.4× every two years) and DRAM capacity (2× every two years). This disparity leads to the memory wall challenge (Gholami et al., 2024), which becomes even more severe when deploying LLMs on memory-limited devices (Zeng et al., 2024; Chen et al., 2024; Hur et al., 2023).
Recently, zeroth-order (ZO) optimization has emerged as a promising memory-efficient training paradigm for LLM fine-tuning, attracting significant attention (Zhang et al., 2024; Liu et al., 2024; Malladi et al., 2023; Zhao et al., 2024). By relying solely on forward passes (i.e., inference) to estimate gradients and update model parameters, ZO bypasses the need for backward propagation and significantly reduces extensive storage requirements for activations, gradients, and optimizer states. As reported in Malladi et al. (2023), fine-tuning LLMs via ZO optimization reduces up to 12$\times$ memory cost. Nevertheless, ZO optimization still exhibits a gap in convergence speed and accuracy compared to the conventional first-order (FO) method (i.e., compute gradient via backpropagation). As shown in Table 1, one can observe that the FO method substantially outperforms ZO method in both accuracy and GPU hours. Though ZO method achieves higher throughput due to its computational simplicity, it requires more than 10× iterations for convergence, dramatically increasing GPU hours. Previous studies typically attribute this gap to the fact that ZO optimization leverages random perturbation for gradient estimation, and thus results in unavoidable estimation error, but without further exploration of other underlying causes (Malladi et al., 2023; Gautam et al., 2024; Zhao et al., 2024).
@@
To bridge this gap, we begin by examining the distinct update patterns shown by ZO and FO methods during LLM fine-tuning. Our analysis reveals a substantial difference in their layer-wise update magnitudes. Specifically, ZO method relies on high-dimensional random search and tends to apply uniform-magnitude updates without considering layer-wise individual characteristics. In contrast, FO method benefits from fine-grained gradient estimation and applies diverse-magnitude updates precisely scaled to the layer-wise individual optimization needs. Motivated by these observations, we are interested in investigating: if we can also provide ZO with diverse-magnitude updates, effectively achieving training acceleration and accuracy improvement.
Drawing on these insights, we propose Divergence-driven Zeroth-Order optimization (DiZO). DiZO conducts divergence-driven layer adaptation by incorporating projections, enabling layer-wise adaptive updates that closely resemble FO approaches. Notably, the projections can be optimized without gradients, ensuring that DiZO retains the appealing backpropagation-free features. Moreover, we validate DiZO on a variety of tasks, including classification and generation, using several LLMs such as RoBERTa-large, the OPT series, and the Llama series. Experimental results show that DiZO substantially decreases training iterations for convergence while maintaining throughput, cutting training GPU hours by up to 48% on diverse datasets. Furthermore, our method can be seamlessly integrated with parameter-efficient tuning techniques like low-rank adaptation (Hu et al., 2021) for additional speedups. DiZO also consistently outperforms the representative ZO baselines and, in some cases, surpasses memory-intensive FO fine-tuning.
@@
  • We introduce DiZO, a novel ZO method using divergence-driven layer adaptation, achieving a learning capacity closely resembling FO while maintaining the throughput benefit of ZO optimization.
@@
Figure 1: Comparison of the training dynamics of ZO and FO methods. The X-axis represents layer names, and the Y-axis represents the distance gap. Although they converge to different stable states, the divergence of the distance gap increases in both FO and ZO methods during training. FO accumulates divergence rapidly through diverse-magnitude updates, while ZO applies uniform-magnitude updates, requiring more iterations for an ideal divergence level. Results are obtained by fine-tuning OPT-2.7B on the SST-2 dataset, focusing on weights in the attention module: K (Key), V (Value), Q (Query), and O (Output projection).
@@ 2.1 Revisiting Zeroth-order Optimization
Recently, ZO optimization has gained significant attention in machine learning (Verma et al., 2023; Dhurandhar et al., 2019; Wang et al., 2022; Gu et al., 2021). Unlike conventional FO optimization, which calculates gradients via backpropagation, ZO optimization estimates gradients using only objective oracles via finite differences (Chen et al., 2023; Liu et al., 2018; Ye et al., 2018). This property can be leveraged for LLM fine-tuning to alleviate the extensive memory costs. Specifically, as ZO only needs two forward passes to obtain the estimated gradients, it avoids computing and storing the most memory-consuming information needed in the conventional FO training, i.e., activations in the forward process, gradients in the backward process, and the optimizer state.
The core idea of ZO optimization is to estimate gradients by applying random perturbations to the weights and computing differences in the objective. For a mini-batch of data $\mathcal{B}$, sampled from a labeled dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{|\mathcal{D}|}$, a model with parameters $\bm{\theta}\in\mathbb{R}^{d}$, where $d$ represents the dimension of the parameter space, and the corresponding loss function $\mathcal{L}(\bm{\theta};\mathcal{B})$. The gradient is estimated as follows:
@@
where $\bm{u}_{i}$ is a random vector with the same dimension as the model weights and is typically drawn from standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ (Malladi et al., 2023; Zhang et al., 2024), or from Gaussian sampling over a unit sphere (Liu et al., 2018; Shamir, 2017), $q$ is the number of objective queries, and $\epsilon>0$ is a small perturbation scalar for smoothing.
Model Type Acc. Memory
#Train
Iter.
GPU
Hours
RoBERTa FO 91.9 9.2 GB 6.6% 12.3%
ZO 90.5 4.5 GB 100.0% 100.0%
OPT-2.7B FO 94.2 45.4 GB 7.5% 16.8%
ZO 90.0 6.8 GB 100.0% 100.0%
@@ 2.2 Layer-wise Divergence Analysis
Drawing insight from the update formula of ZO optimization, we notice that ZO method applies uniform-magnitude updates across layers, e.g., the L2-norm of the updates is about the same for all layers in one iteration (see Appendix E for proof). This fact may be the root of the inferior performance of ZO optimization. To measure how the divergence of update magnitude affects the convergence speed and accuracy, we investigate the training dynamics of ZO and FO methods, respectively.
Analysis indicator. To quantify the effect of updates, we adapt the layer-wise L2-norm distance gap between the weights of the pre-trained and the fine-tuned model as an indicator. The layer-wise L2-norm distance gap is defined as:
@@
Analysis result. Figure 1 compares the training dynamics of FO and ZO methods. Regardless of whether ZO or FO is used, the divergence of distance gap among layers grows during training, i.e., the line of distance gap gradually ’bends’. This pattern implies that different layers benefit from maintaining diverse distance gaps with the pre-trained model. However, FO and ZO diverge in how the distance gap divergence is accumulated. FO employs fine-grained gradient estimations, resulting in diverse-magnitude updates (FO updates in Figure 1). Therefore, it can rapidly reach the desired layer-wise distance gap in only a few iterations. In contrast, ZO relies on random search in high-dimensional parameter space and generates uniform-magnitude updates (ZO updates in Figure 1), resulting in thousands more iterations required for accumulating a meaningful layer-wise distance gap.
With the above findings, we suspect the inferior performance of ZO stems from its inability to deliver layer-wise adaptive updates, a challenge that arises from its reliance on random perturbations for gradient estimation.
@@
Require: parameter of $t$-th iteration $\bm{\theta_{t}}$ and pre-trained model $\bm{\theta_{0}}$, loss function $\mathcal{L}$, step budget $T$, perturbation scalar $\epsilon$, mini-batch data $\mathcal{B}_{t}$, learning rate $\eta$, projection at $t$-th iteration $\bm{\gamma_{t}}=\{\gamma_{t}^{i}\}_{i=1}^{L}$
for $t=1$ to $T$ do
       $\nabla\mathcal{L}=\text{{GradEst}}(\bm{\theta_{t}},\epsilon,\mathcal{B}_{t})$$\bm{\theta_{t}}=\bm{\theta_{t-1}}-\eta\nabla\mathcal{L}$$\bm{\gamma}^{*}_{t}=\arg\min_{\bm{\gamma}_{t}}\mathcal{L}(\bm{\theta}_{0}+\frac{\bm{\gamma}_{t}}{\|\Delta\bm{\theta}_{t}\|}\Delta\bm{\theta}_{t};\mathcal{B}_{t})$$\bm{\theta}_{t}=\text{{ApplyProjection}}(\bm{\theta}_{t},\bm{\theta}_{0},\bm{\gamma}_{t}^{*})$
end for
Subroutine GradEst($\bm{\theta}$, $\epsilon$, $\mathcal{B}$):
       Sample: $u_{1},\dots,u_{q}\backsim\mathcal{N}(0,\mathbf{I})$Query: $y_{i}=\mathcal{L}(\bm{\theta}+\epsilon u_{i};\mathcal{B})-\mathcal{L}(\bm{\theta}-\epsilon u_{i};\mathcal{B})$Estimator: $\nabla\mathcal{L}=\frac{q}{2\epsilon}\sum_{i=1}^{q}y_{i}u_{i}$return $\nabla\mathcal{L}$
return
Subroutine ApplyProjection($\bm{\theta}_{t}$, $\bm{\theta}_{0}$, $\bm{\gamma}_{t}$):
       for $\ell=1,2,\dots,L$ do
             // Project $l$-th layer $\bm{\theta}_{t}^{(\ell)}=\bm{\theta}_{0}^{(\ell)}+\frac{\gamma^{(\ell)}_{t}}{\|\Delta\bm{\theta}_{t}^{(\ell)}\|}\Delta\bm{\theta}_{t}^{(\ell)}$
       end for
      return $\bm{\theta}_{t}$
return
@@ 3 Methodology
We find that ZO applies uniform-magnitude updates for all layers, which could be the root of its inferior performance in accuracy and convergence speed. Consequently, we introduce a variant of ZO optimization which performs divergence-driven layer adaptation, thereby providing diverse-magnitude updates to enhance the overall learning capacity.
@@ 3.1 Design of the Divergence-driven Layer Adaptation
To provide layer-wise adaptive updates for ZO optimization, we apply projections to the updates of different layers, generating updates with diverse magnitudes. The pseudocode for the proposed method is shown in Algorithm 1.
@@
where we get the new $\bm{\theta}_{t}$ after projection, and then we use the projected one for the following fine-tuning. When the value of $\bm{\gamma_{t}}$ is larger than $\|\Delta\bm{\theta_{t}}\|$, enlarges the distance gap between the fine-tuned model and the pre-trained model , and vice versa.
@@ 3.2 How to Learn the Projection?
Although promising, finding the ideal projection (defined in Eq. (4)) remains challenging due to the high complexity of the objective. A straightforward solution is to also perform backpropagation for gradient computation and optimize the projection accordingly (FO-based method). For example, we use Adam optimizer to directly update $\bm{\gamma}_{t}$. The results are shown in Table 2, one can observe that it significantly reduces 67.7% of the iteration and 58.5% of the training GPU hours, and increases by 3.4% in accuracy. These results underscore the effectiveness of incorporating our proposed divergence-driven layer adaptation.
However, searching projection with the FO method makes DiZO only partially gradient-free. Specifically, while the model weights are updated via ZO, the per-layer projection parameter $\gamma_{t}^{(\ell)}$ is updated via FO, which still requires the backward pass and storing memory-intensive activation. The only memory saving, compared to the vanilla FO fine-tuning, is the optimizer state. As a result, relying on FO to find the ideal projection, though it achieves faster convergence speed and better accuracy in ZO optimization, offers limited overall benefit. It is worth noting that the peak memory usage during training of the FO-based DiZO is similar to that of low-rank adaptation (LoRA) (Hu et al., 2021).
Is the projection-based method for enhancing layer-wise divergence in ZO a failed idea that seems promising at first glance but is actually not after deliberation? Fortunately, the answer is no. We develop a ZO projection learning algorithm, which retains the memory-efficient advantages and also achieves training acceleration and accuracy enhancement.

### v3

@@ Abstract
Large language models (LLMs) excel across various tasks, but standard first-order (FO) fine-tuning demands considerable memory, significantly limiting real-world deployment. Recently, zeroth-order (ZO) optimization stood out as a promising memory-efficient training paradigm, avoiding backward passes and relying solely on forward passes for gradient estimation, making it attractive for resource-constrained scenarios. However, ZO method lags far behind FO method in both convergence speed and accuracy. To bridge the gap, we introduce a novel layer-wise divergence analysis that uncovers the distinct update pattern of FO and ZO optimization. Aiming to resemble the learning capacity of FO method from the findings, we propose Divergence-driven Zeroth-Order (DiZO) optimization. DiZO conducts divergence-driven layer adaptation by incorporating projections to ZO updates, generating diverse-magnitude updates precisely scaled to layer-wise individual optimization needs. Our results demonstrate that DiZO significantly reduces the needed iterations for convergence without sacrificing throughput, cutting training GPU hours by up to 48% on various datasets. Moreover, DiZO consistently outperforms the representative ZO baselines in fine-tuning RoBERTa-large, OPT-series, and Llama-series on downstream tasks and, in some cases, even surpasses memory-intensive FO fine-tuning. Our code is released at https://github.com/Skilteee/DiZO.
@@ 1 Introduction
Fine-tuning large language models (LLMs) via backpropagation achieves strong performance across many NLP tasks (yang2019end, ; liu2019roberta, ; talmor2018commonsenseqa, ; chowdhery2023palm, ; zheng2020end, ), but their large parameter counts create substantial memory burdens, limiting downstream applicability. Following neural scaling laws (hoffmann2022empirical, ; kaplan2020scaling, ), next-generation LLMs grow rapidly, e.g., model sizes increase 410× every two years, far outpacing DRAM bandwidth (1.4×) and capacity (2×) growth. This imbalance leads to the memory wall(gholami2024ai, ), a growing challenge especially for deployment on memory-limited devices(zeng2024flightllm, ; chen2024understanding, ; tan2025perturbation, ).
[…]
Model Type Acc. Memory
GPU
Hours
RoBERTa FO 91.9 9.2 GB 12.3%
ZO 90.5 4.5 GB 100.0%
OPT-2.7B FO 94.2 45.4 GB 16.8%
ZO 90.0 6.8 GB 100.0%
@@
To further uncover the fundamental causes of this gap, we begin by analyzing the distinct update patterns shown by FO and ZO methods during LLM fine-tuning. Interestingly, our analysis reveals a substantial difference in the magnitude of weight updates between layers. Specifically, FO methods benefit from fine-grained gradient estimation and enable diverse-magnitude updates precisely scaled to the layer-wise individual optimization needs. In contrast, ZO method tends to behave with uniform-magnitude updates without considering layer-wise individual characteristics. This is attributed to the nature of ZO that relies on high-dimensional random search and leverages random perturbation for gradient estimation. Based on this, we conjecture that the compromised performance of ZO stems from its limited capability in achieving diverse-magnitude updates. This naturally raises the question: if we could enable ZO to achieve the desired diverse-magnitude updates, could we effectively achieve training acceleration and accuracy improvement?
To validate our conjecture and fill the performance gap, we innovatively propose Divergence-driven Zeroth-Order optimization (DiZO), which performs divergence-driven layer adaptation via anchor-based learnable projections, enabling principled adaptive updates that resemble FO methods. Specifically, DiZO guides updates along geometrically constrained directions by learning target distances from an anchor point (e.g., the pre-trained model). We also design a ZO-based method for projection learning that ensures the entire training process is memory-efficient. We extensively evaluate DiZO on a range of tasks, including classification and generation, using models such as RoBERTa-large, the OPT series, and the Llama series. Results show that DiZO significantly reduces training iterations and cuts GPU hours by up to 48% without sacrificing throughput. DiZO also integrates seamlessly with parameter-efficient tuning methods like LoRA (hu2021lora, ), and consistently outperforms ZO baselines, sometimes even surpassing memory-intensive FO fine-tuning. Finally, we comprehensively analyze several potential alternatives and validate the necessity and effectiveness of our approach.
@@
  • We introduce DiZO, a novel ZO method using divergence-driven layer adaptation, achieving a learning capacity closely resembling FO while maintaining the throughput benefit.
@@
  • We also provide comprehensive analysis and discussions on overheads, convergence guarantee, and potential alternatives, which further strengthen the efficiency, feasibility, and necessity of our proposed approach.
Figure 1: Comparison of the training dynamics of ZO and FO methods. For the upper subfigure, $W_{K},W_{V},W_{Q},W_{O}$ indicate the corresponding weight matrix in the attention module.
@@ 2.1 Revisiting Zeroth-order Optimization
Recently, ZO optimization has gained significant attention in machine learning (verma2023certified, ; dhurandhar2019model, ; wang2022zarts, ; gu2021efficient, ). Unlike conventional FO optimization, which calculates gradients via backpropagation, ZO optimization estimates gradients using only objective oracles via finite differences (chen2023deepzero, ; liu2018zeroth, ; ye2018hessian, ). This property can be leveraged for LLM fine-tuning to alleviate the extensive memory costs. Specifically, as ZO only needs two forward passes to obtain the estimated gradients, it avoids computing and storing the most memory-consuming information needed in the conventional FO training, i.e., activations in the forward process, gradients in the backward process, and the optimizer state.
The core idea of ZO optimization is to estimate gradients by applying random perturbations to the weights and computing differences in the objective. For a mini-batch of data $\mathcal{B}$, sampled from a labeled dataset $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{|\mathcal{D}|}$, a model with parameters $\bm{\theta}\in\mathbb{R}^{d}$, where $d$ represents the dimension of the parameter space, and the corresponding loss function $\mathcal{L}(\bm{\theta};\mathcal{B})$. The gradient is estimated as:
@@
where $\bm{u}_{i}\sim\mathcal{N}(0,\mathbf{I})$ is a random perturbation typically drawn from standard Gaussian distribution, $q$ is the number of queries, and $\epsilon>0$ is a small perturbation scalar for smoothing.
@@ 2.2 Layer-wise Divergence Analysis
ZO optimization applies uniform-magnitude updates across layers, with similar L2-norms per iteration (see Appendix F). This uniformity may underlie its weaker performance. To explore how update divergence impacts convergence and accuracy, we analyze the training dynamics of ZO and FO methods.
Analysis indicator. To quantify the effect of updates, we adapt the layer-wise L2-norm distance gap between the weights of the pre-trained and the weights of fine-tuned model as an indicator. The layer-wise L2-norm distance gap is defined as:
@@
Analysis result. Figure 1 compares the training dynamics of FO and ZO optimization methods, focusing on how each shapes the layer-wise distance gap between model parameters and the pre-trained initialization. Both methods show increasing divergence among layers over time, as illustrated by the twisting lines in the upper sub-figure, indicating that layers benefit from different levels of deviation from the pre-trained model. However, FO and ZO differ in how this divergence accumulates. FO uses fine-grained gradients to produce diverse-magnitude updates, quickly establishing meaningful distance gaps within a few iterations. In contrast, ZO performs random search with uniform-magnitude updates, requiring thousands of iterations to reach similar divergence. Additionally, FO-based methods show fast but converging growth in distance gaps, while ZO-based methods exhibit linear, unconstrained growth. This continued expansion under ZO may reflect a lack of effective constraint, contributing to accuracy drops in later stages and resulting in suboptimal performance.
1Require: parameter of $t$-th iteration $\bm{\theta_{t}}$ and pre-trained model $\bm{\theta_{0}}$, loss function $\mathcal{L}$, step budget $T$, perturbation scalar $\epsilon$, mini-batch data $\mathcal{B}_{t}$, learning rate $\eta$, projection at $t$-th iteration $\bm{\gamma_{t}}=\{\gamma_{t}^{i}\}_{i=1}^{L}$, projection update interval $\kappa$
2for $t=1$ to $T$ do
3  $\nabla\mathcal{L}=\text{{GradEst}}(\bm{\theta_{t}},\epsilon,\mathcal{B}_{t})$;
4
5 $\bm{\theta_{t}}=\bm{\theta_{t-1}}-\eta\nabla\mathcal{L}$;
6
7 if $t\bmod\kappa=0$ then
8     $\bm{\gamma}^{*}_{t}=\arg\min_{\bm{\gamma}_{t}}\mathcal{L}(\bm{\theta}_{0}+\frac{\bm{\gamma}_{t}}{\|\Delta\bm{\theta}_{t}\|}\Delta\bm{\theta}_{t};\mathcal{B}_{t})$;
9
10    $\bm{\theta}_{t}=\text{{ApplyProjection}}(\bm{\theta}_{t},\bm{\theta}_{0},\bm{\gamma}_{t}^{*})$;
11
12  end if
13
14 end for
15
16 Subroutine GradEst($\bm{\theta}$, $\epsilon$, $\mathcal{B}$):
17  Sample: $u_{1},\dots,u_{q}\backsim\mathcal{N}(0,\mathbf{I})$;
18
19 Query: $y_{i}=\mathcal{L}(\bm{\theta}+\epsilon u_{i};\mathcal{B})-\mathcal{L}(\bm{\theta}-\epsilon u_{i};\mathcal{B})$;
20
21 Estimator: $\nabla\mathcal{L}=\frac{q}{2\epsilon}\sum_{i=1}^{q}y_{i}u_{i}$;
22
23 return $\nabla\mathcal{L}$;
24
25 return
26
27 Subroutine ApplyProjection($\bm{\theta}_{t}$, $\bm{\theta}_{0}$, $\bm{\gamma}_{t}$):
28  for $\ell=1,2,\dots,L$ do
29
    // Project $l$-th layer
30     $\bm{\theta}_{t}^{(\ell)}=\bm{\theta}_{0}^{(\ell)}+\frac{\gamma^{(\ell)}_{t}}{\|\Delta\bm{\theta}_{t}^{(\ell)}\|}\Delta\bm{\theta}_{t}^{(\ell)}$;
31
32  end for
33 return $\bm{\theta}_{t}$;
34
35 return
@@ 3.1 Design of the Divergence-driven Layer Adaptation
To provide layer-wise adaptive updates for ZO optimization, we propose applying anchor-based learnable projections to the updates of different layers. The pseudocode for the method is shown in Algorithm 1.
@@
where we get the new $\bm{\theta}_{t}$ after projection, and then we use the projected one for the following fine-tuning. When the value of $\bm{\gamma_{t}}$ is larger than $\|\Delta\bm{\theta_{t}}\|$, it enlarges the distance gap between the fine-tuned model and the pre-trained model, and vice versa.
@@ 3.2 How to Learn the Projection?
Although promising, finding the ideal projection (defined in Eq. (4)) remains challenging due to the high complexity of the objective. A straightforward solution is to also perform backpropagation for gradient computation and optimize the projection accordingly (FO-based method). For example, we use Adam optimizer to directly update $\bm{\gamma}_{t}$. The results are shown in Table 2, which significantly reduces 67.7% of the iterations and 58.5% of GPU hours, and increases by 3.4% in accuracy.
@@
However, searching projection with the FO method makes DiZO only partially gradient-free. Specifically, while the model weights are updated via ZO, the per-layer projection parameter $\gamma_{t}^{(\ell)}$ is updated via FO, which still requires the backward pass and storing memory-intensive activation. The only memory saving, compared to the vanilla FO fine-tuning, is the optimizer state. As a result, relying on FO to find the ideal projection, though it achieves faster convergence speed and better accuracy in ZO optimization, offers limited overall benefit. It is worth noting that the peak memory usage during training of the FO-based DiZO is similar to that of low-rank adaptation (LoRA) (hu2021lora, ).